OpenStudy (xxstudyingstuffxx):

Simplify 2m^2(-3m + 4m^2) - (3m)^3

OpenStudy (michele_laino):

hint: we have this: \[2{m^2} \cdot \left( { - 3m} \right) = - 6{m^3}\] \[2{m^2} \cdot 4{m^2} = 8{m^4}\]

OpenStudy (michele_laino):

now, please apply the distributive property of multiplicationn over addition: \(a(b+c)=ab+ac\)

OpenStudy (michele_laino):

multiplication*

OpenStudy (xxstudyingstuffxx):

Thanks!

OpenStudy (michele_laino):

:)